Mark​ Gershon, owner of a musical instrument​ distributorship, thinks that demand for guitars may be related to the number of television appearances by the popular group Maroon 5 during the previous month. Gershon has collected the data shown in the following​ table:

                                                                                                          

Maroon 5 TV Appearances

33

55

77

55

77

55

Demand for Guitars

44

66

88

66

1111

66

This exercise contains only parts​ b, c, and d.

​b) Using the​ least-squares regression​ method, the equation for forecasting is ​(round your responses to four decimal

​places)​:

Y​ =

nothing

​+

nothingx

​c) The estimate for guitar sales if Maroon 5 performed on TV

99

times​ =

nothing

sales ​(round your response to two decimal​ places).​d) The correlation coefficient

​(r​)

for this model​ =

nothing

​(round your response to four decimal​ places).The coefficient of determination

​(r squaredr2​)

for this model​ =

nothing

​(round your response to four decimal​ places).The percentage of variation in sales that can be explained by TV appearances​ =

nothing​%

​(round your response to two decimal​ places).
Answer #1

b))HZ Data Analysis ? X к 1 Maroon 5 appearances Demand for guitars ок | Cancel Analysis Tools Histogram Moving Average Random N

. Get Refresh 2 Fliter Text to Data Stocks Geography - All E dit Links Get & Transform Data Queries & Connections RegressionPU B C D E F G H I J 1 SUMMARY OUTPUT 2 3 Regression Statistics 4 Multiple R 0.903542463 5 R Square 0.816388983 6 Adjusted R

The equation is

Y = -0.8529 + 1.4412X

c)

When X = 9

Y = -0.8529 +1.4412*9 = 12.12 guitars

d)

Percentage of variation explained is given by R square value

Percentage of variation explained = 81.64 %